# AdaptInfer: Adaptive Token Pruning for Vision–Language Model Inference with Dynamical Text Guidance

### v1

@@ Appendix D Visualization
The experiment is carried out on the TextVQA (Singh et al. 2019) benchmark, from which we sample a representative set of examples for display. As the layers deepen, AdaptInfer rapidly suppresses tokens associated with semantically irrelevant background areas while consistently preserving those aligned with question-critical cues. The visualization is shown in Figure A.7.
Appendix E Performance Analysis
Figure A.8: Examples where AdaptInfer outperforms SparseVLM. AdaptInfer performs better when text prompts are rich in descriptive modifiers.
We provide further analysis of the advantages of our dynamic text guidance method, AdaptInfer, compared with the static text guidance solution, SparseVLM (Zhang et al. 2025). We extract some VQA examples from the TextVQA dataset (Singh et al. 2019), and the answers of both methods are presented in Figure A.8.
@@ Appendix E Performance Analysis
Appendix F Experiment Setup
@@
Implement Detail
In the paper, we evaluated three vision–language models, including LLaVA-1.5-7B, LLaVA-1.5-13B, and InternVL-Chat-7B (Chen et al. 2024a) in total. To stay within GPU-memory limits, we ran the LLava-7B variant on one GPU, the LLava-13B variant on three GPUs, and InternVL-Chat-7B on two GPUs (InternVL-Chat-7B contains a 6B-parameter vision encoder). All training and inference are performed in torch.float16 precision to further conserve memory usage. All the results of AdaptInfer reported in the context are the average of five runs.
Randomness Report

### v2

@@ Appendix G Visualization
The experiment is carried out on the TextVQA (Singh et al., 2019) benchmark, from which we sample a representative set of examples for display. As the layers deepen, AdaptInfer rapidly suppresses tokens associated with semantically irrelevant background areas while consistently preserving those aligned with question-critical cues. The visualization is shown in Figure 6.
Appendix H Performance Analysis
Figure 7: Examples where AdaptInfer outperforms SparseVLM. AdaptInfer performs better when text prompts are rich in descriptive modifiers.
We provide further analysis of the advantages of our dynamic text guidance method, AdaptInfer, compared with the static text guidance solution, SparseVLM (Zhang et al., 2025). We extract some VQA examples from the TextVQA dataset (Singh et al., 2019), and the answers of both methods are presented in Figure 7.
@@ Appendix H Performance Analysis
Appendix I Experiment Setup
@@ I.2 Software Configuration
I.3 General Implementation Details
In our evaluation of AdaptInfer, the pruning ratios at each pruning location are the hyperparameters need to be selected as well. For these hyperparameters, we follow the same rank-based sparsification-level adaptation method which SparseVLM (Zhang et al., 2025) introduced.
In the paper, we evaluated four vision–language models in total, including LLaVA-1.5-7B, LLaVA-1.5-13B, InternVL-Chat-7B (Chen et al., 2024a) and Qwen2-VL-2B. To stay within GPU-memory limits, we ran the LLava-7B variant and Qwen2-VL-2B on one GPU, the LLava-13B variant on three GPUs, and InternVL-Chat-7B on two GPUs (InternVL-Chat-7B contains a 6B-parameter vision encoder). All training and inference are performed in torch.float16 or torch.bfloat16 precision to further conserve memory usage. All the results of AdaptInfer reported in the context are the average of five runs.
I.4 Implementation Details of AdaptInfer on Qwen2-VL-2B
Qwen2-VL-2B employs a dynamic number of vision tokens, ranging from 4 to 16,384 per image. To ensure stable performance and prevent GPU memory overflow, we constrain the vision token count within a bounded range, setting the minimum to 256 and the maximum to 1,280. For video-based multimodal evaluation, we follow FastV (Chen et al., 2025a) and SparseVLM (Zhang et al., 2025) and use the first 1,000 test samples from each benchmark to reduce evaluation time. The frame rate for all videos is fixed at 4 FPS.
I.5 Compatibility with FlashAttention
During token ranking, we need t2t and t2v attentions to calculate the token importance scores. However, the original FlashAttention (Dao et al., 2022) does not allowed attention extraction. Fortunately, this issue has already been addressed by SparseVLM (Zhang et al., 2025), which provides a lightweight dual-pass solution that remains fully compatible with FlashAttention. Our method directly follows the same strategy.
Notably, because this solution can only get the mean t2t and t2v attention scores (averaged by multi-attention heads), we need to adjust our eqaution 2 and 3 in the main paper into:
$\mathbf{s}=({\sum_{i=1}^{T}\mathbf{A}^{(mean)}_{t2t}}[i:])^{\top}\cdot\mathbf{A}^{(mean)}_{\text{t2v}}\in\mathbb{R}^{V}.$ (11)
I.6 Randomness Report